OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x2 find g(f(x))

OpenStudy (anonymous):

So, g(f(x)) = (3x+2)^2 = (3x+2)(3x+2) = 9x^2 + 12x + 4

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

f(x) => x = 3x +2 g (f(x)) => x^2= (3x+2) ^2 = 9x^2 + 12x +4

OpenStudy (anonymous):

np :)